OpenStudy (anonymous):

Simplify the radical. the square root of 90 A. 9 [10] B. 10 [9] C. 3 [10] D. 10 [3] Simplify the radical. the square root of 90 A. 9 [10] B. 10 [9] C. 3 [10] D. 10 [3] @Mathematics

OpenStudy (asnaseer):

\[\sqrt{90}=\sqrt{9*10}=\sqrt{9}*\sqrt{10}=3\sqrt{10}\]

OpenStudy (asnaseer):

I hope the way the answer was discovered made sense to you.